My question has to be solved using elimination. The two equations in the question are:
5x+2y=18 and,
3x+2y=10.
How can I solve this and what is the correct answer?
Thank you!
Amber Robinson

Answer by checkley75(3666)

5X+2Y=18
2Y=-5X+18 NOW SUSTITUE (-5X+18) FOR 2Y & SOLVE FOR X IN THE SECOND
3X+(-5X+18)=10
3X-5X+18=10
-2X=10-18
-2X=-8
X=-8/-2
X=4 ANSWER.
5*4+2Y=18
20+2Y=18
2Y=18-20
2Y=-2
Y=-2/2
Y=-1 ANSWER.
PROOF
3*4+2*-1=10
12-2=10
10=10